so the question says find the measure of all 4 angles of the quadrilateral. it gives me 4 angles one of which is 3x and 4x+12. How do I find the angles?
Answer by rwm(914)

A quad's angles total 360.
a quad has 4 angles. Two sets of two which are equal
so we have 2*3x+2*(4x+12)=360
6x+8x+24=360
14x=360-24
14x=336
7x=118
x=24
one angle is 3x=3*24=72
the other is 4x+12=4*24+12=96+12=108
so we have 72, 108, 72, 108